OpenStudy (anonymous):

The number of cars at a dealership dropped from 64 to 48 after a weekend sale. What is the percent of decrease in the number of cars?

OpenStudy (anonymous):

the question boils down to, 48 is what percent of 64? 64/48 = x 100/x = answer

OpenStudy (anonymous):

On My Calculator Im Getting 133.333333333333333

OpenStudy (anonymous):

not quite sure what you did, but you are on the right track. it's 100 divided by whatever 64/48 is

OpenStudy (anonymous):

Ohhh Okay,Thank You Guys.

OpenStudy (anonymous):

the answer isn't 25 either

OpenStudy (anonymous):

I Know But Im Saying Thank You For Helping Me.

OpenStudy (anonymous):

alright, just checking lol and you're welcome!

OpenStudy (anonymous):

Okay So 64 - 48 Is 16 Which Means 16 * 100 =1,600

OpenStudy (anonymous):

ah nevermind i feel bad now, the answer is 25%

OpenStudy (anonymous):

it's the percent decrease, not the the difference in the numbers

OpenStudy (anonymous):

and it's not * 100, /100, you want decimals